# A signal separation view of classification

### v1

@@ 4 Background
where $\Phi_{n}$ is the kernel defined in (3.3). The localization property (3.5), used with $\lceil S/2\rceil$ in place of $S$ implies that
@@ 4.1 Measures
In the case of signal separation, we have seen that if the minimal amplitude for a certain point source is sufficiently small, we may not be able to detect that point source. Likewise, if the measure $\mu$ is too small on parts of $\mathbb{X}$}, we may not be able to detect those parts. For this reason, we make some assumptions on the measure $\mu$ as in [7]. The first property, detectability, determines the rate of growth of the measure locally around each point in the support. The second property, fine-structure, relates the measure to the classification problem by equipping the support with some well-separated (except maybe for some subset of relatively small measure) partition which may correspond to some different class labels in the data.
@@
Supposing that $\mu=\sum_{k=1}^{K}a_{k}\delta_{\omega_{k}}$ (as in the signal separation problem), then we see that $\mu$ is detectable with $\alpha=0$, $\kappa_{1}=\max_{k}|a_{k}|$, $\kappa_{2}=\min_{k}|a_{k}|$. It has fine structure in the classical sense whenever $\eta<\min_{j\neq k}|\omega_{j}-\omega_{k}|$. In this sense, the theory presented in this paper is a generalization of results for signal separation in this regime. ∎
@@ 4.2 F-score
We will give results on the theoretical performance of our measure estimating procedure by giving an asymptotic result involving the so-called F-score. The F-score for binary classification (true/false) problems is a measure of classification accuracy taking the form of the harmonic mean between precision and recall. In a predictive model, precision is defined as the fraction of true positive outputs over all the positive outputs of the model. Recall is the fraction of true positive outputs over all the actual positives. In a multi-class problem, we extend this definition as follows (cf. [26]). If $\{C_{1},\dots,C_{N}\}$ is a partition of $\{x_{j}\}_{j=1}^{M}$ indicating the predicted output labels of a model and $\{L_{1},\dots,L_{K}\}$ is the ground-truth partition of the data, then one can define the precision of $C_{j}$ against the true label $L_{k}$ by $|C_{j}\cap L_{k}|/|C_{j}|$ and the corresponding recall by $|C_{j}\cap L_{k}|/|L_{k}|$. Taking the maximum of the harmonic means of the precisions and recalls with respect to all the ground truth labels leads to
@@ 5 Main results
Our first theorem gives us bounds on how well $\mathcal{G}_{n}(\Theta)$ approximates the entire support $\mathbb{X}$ with the detectability assumption.
@@
Let $\mu$ be detectable and suppose $M\gtrsim n^{\alpha}\log(n)$. Let $\{x_{1},x_{2},\dots,x_{M}\}$ be independent samples from $\mu$. There exists a constant $C>0$ such that if $\Theta<C<1$, then there exists $r(\Theta)\sim\Theta^{-1/(S-\alpha)}$ such that with probability at least $1-c_{1}/M^{c_{2}}$ we have
@@
If $\mathcal{C}=\{z_{1},\cdots,z_{M}\}$ is a random sample from $\mu$, $n\geq 1$ and $M\gtrsim n^{\alpha}\log n$, then it can be shown (cf. [17, Lemma 7.1]) that for any point $x\in{\mathbb{X}}$, there exists some $z\in\mathcal{C}$ such that $\rho(x,z)\leq 1/n$. Hence, the Hausdorff distance between $\mathbb{X}$} and $\mathcal{C}$ is $\leq 1/n$. If $\mu$ has a fine structure in the classical sense, and $n\gtrsim\eta^{-1}$, then this implies that a correct clustering of $\mathcal{C}$ would give rise to a correct classification of every point in $\mathbb{X}$}. This justifies our decision to construct the algorithm in Section 6 to classify only the points in $\mathcal{C}$. On the other hand, the use of the localized kernel as in the theorems above guide us about the choice of the points at which to query the label. ∎
@@
Figure 3: Demonstration of the support estimation set $\mathcal{G}_{32}(0.15)$ (yellow) applied to a simple two-moons data set from [11] (blue and red). By querying one point from each component of the support estimation set and extending the label to the other points in the same component, we can classify the entire data set with 100% accuracy.
@@ 6.1 Algorithm description
MASC will collect all points which do not yet have predicted labels (line 17), and apply the nearest-neighbors approach as describe above to each of these points (lines 19 and 20). At this point, every element in the data set will have a predicted label, so the algorithm will return the list of labels (line 21).
@@ 6.2 Comparison with CAC and SCALe
In [7], a similar theoretical approach to this paper except on the Euclidean space was developed and an algorithm we will call “Cautious Active Clustering” (CAC) was introduced. MASC and CAC are both multiscale algorithms using $\mathcal{G}_{n}(\Theta)$ to threshold the data set, then constructing graphs to query points and extend labels. The main difference between the algorithms is the following. In CAC, $\eta,\Theta$ are considered hyperparameters while $n$ is incremented, whereas in MASC, $n,\Theta$ are considered hyperparameters while $\eta$ is incremented. This adjustment serves three purposes:
@@
SCALe, as introduced in [19] is an even more similar algorithm to MASC. The main difference between MASC and SCALe is the final step, where in the present method we use a nearest-neighbors approach to extend labels to uncertain points while in SCALe the choice was to use a function approximation technique developed in [20]. Both methods have their pros and cons. Compared to SCALe, the nearest-neighbors approach of MASC:
@@
    Our algorithm works in arbitrary metric spaces, without requiring a summability kernel as in SCALe.
@@ 7 Numerical examples
In this section, we look at the performance of the MASC algorithm applied to 1) a synthetic data set with overlapping class supports (Section 7.1), 2) a document data set (Section 7.2), and 3) two different hyperspectral imaging data sets: Salinas (Section 7.3) and Indian Pines (Section 7.4). On the Hyperspectral data sets, we compare our method with two other algorithms for active learning: LAND and LEND (Section 7.5).
For hyperparameter selection on our model as well as the comparisons, we have not done any validation but rather optimized the hyperparameters for each model on the data itself. So the results should be interpreted as being near-best-possible for the models applied to the data sets in question rather than a demonstration of generalization capabilities. While this approach is non-traditional for unsupervised/supervised learning, it has been done for other active learning research ([31], for example) so we have elected to follow the same procedure in this paper. Further, an exhaustive grid search was not conducted but rather local minima among grid values were selected for each hyperparameter. For MASC, we looked at $n$ in powers of $2$ and $\overline{k}$ values in multiples of $5$. For LAND we looked at $K,t$ at increments of $10$, and with LEND we used the same parameters from LAND and looked at integer $J$ values and $\alpha$ values in increments of $0.1$. For $\Theta$, we tried values less rigorously, meaning that better $\Theta$ values may exist than the ones chosen. Due to the nature of the algorithm, increasing $\Theta$ will increase the number of samples that the nearest-neighbors approach has to estimate, while reducing the number of labeled neighbors it has to do so. However, increasing $\Theta$ can also reduce the number of queries used, sometimes without deterioration in accuracy. So there may be some tradeoff, but we generally see the best results when $\Theta$ is chosen to threshold a small portion of the initial data (outlier removal). In Table 2, we summarize the choice of parameters for each of the data sets in the subsequent sections.
@@ 7.2 Document data
This numerical example uses the document data set provided by Jensen Baxter through Kaggle [2]. The data set contains 1000 documents total, 100 each belonging to a particular category from: business, entertainment, food, graphics, historical, medical, politics, space, sport, and technology. For prepossessing we run the data through the Python sklearn package’s TfidfVectorizer function to convert the documents into vectors of length 1684. Then we implement MASC.
@@
Figure 5: This figure illustrates the classification process undergone MASC on the document data set at two points. On the left, we see the classification of points before the $\overline{k}$-nearest neighbors extension. On the right, we see the result after $\overline{k}$-nearest neighbors extension.
@@ 7.3 Salinas hyperspectral data
This numerical example is done on a subset of the Salinas hyperspectral image data set from [12]. Our subset of the Salinas data set consists of 20034 data vectors of length 204 belonging to 10 classes of the 16 original classes. Specifically, we took half of the data points at random from each of the first 10 classes of the original data set. For preprocessing we ran PCA and kept the first 50 components. Then we implemented MASC.
@@
Figure 8: This figure illustrates the classification process undergone by MASC at two points on the Salinas hyperspectral data set. On the left, we see the classification of points before the $\overline{k}$-nearest neighbors extension. On the right, we see the result after $\overline{k}$-nearest neighbors extension.
@@ 7.4 Indian Pines hyperspectral data
This numerical example is done on a 5-class subset of the Indian Pines hyperspectral image data set from [12]. Our subset of the Indian Pines data set consists of 5971 data vectors of length 200 belonging to classes number 2,6,11,14,16 of the 16 original classes. For preprocessing we normalized each vector. Then we implement MASC.
@@
Figure 11: This figure illustrates the classification process undergone by MASC at two points on the Salinas hyperspectral data set. On the left, we see the classification of points before the $\overline{k}$-nearest neighbors extension. On the right, we see the result after $\overline{k}$-nearest neighbors extension.
@@ 7.5 Comparison with LAND and LEND
We compare our method with the LAND [15] algorithm and its boosted variant, LEND [31]. In Figure 14, we see the resulting accuracy that each algorithm achieves on both Salinas and Indian Pines for various query budgets. On the left, we observe that our method achieves a comparable accuracy to both LAND and LEND at around 50 queries, then gradually surpasses the accuracy of LAND as the number of queries surpasses around 200. On the right, our method achieves a lower accuracy for a small number of queries, but then outperforms both LAND and LEND after the budget exceeds about 60 queries.
@@ 8.1 Measure support estimation
In order to prove this lemma, we first recall a consequence of the Bernstein inequality for trigonometric polynomials ([21], Chapter III, Section 3, Theorems 1 and Lemma 5).
@@ 8.2 Discretization
Information and a derivation for the Bernstein concentration inequality are standard among many texts in probability; we list [3, Section 2.1, 2.7] as a reference. It is instinctive to use Proposition 8.1 with $X_{j}=\Psi_{n}(x,x_{j})$. This would yield the desired bound for any value of $x\in\mathbb{M}$. However, to get an estimate on the supremum norm of the difference $F_{n}-\sigma_{n}$, we need to find an appropriate net for $\mathbb{M}$ (and estimate its size) so that the point where this supremum is attained is within the right ball around one of the points of the net. Usually, this is done via a Bernstein inequality for the gradients of the objects involved. In the absence of any differentiability structure on $\mathbb{M}$, we need a more elaborate argument.
The following proposition is a consequence of [17, Theorem 7.2], and asserts the existence of a partition of $\mathbb{X}$ satisfying properties which will be helpful to proving our main results.
@@ 9 Conclusions
We also give some empirical analysis for the performance of our new algorithm MASC, which was originally introduced in a varied form in [19]. The key focus of the algorithm is on querying high-information points whose labels can be extended to others belonging to the same class with high probability. This is done in a multiscale manner, with the intention to be applied to data sets where the minimal separation between supports of the measures for different classes may be unknown or even zero. We applied MASC to a document data set as well as two hyperspectral data sets, namely subsets of the Indian Pines and Salinas hyperspectral imaging data sets. In the process of these experiments, we demonstrate empirically that MASC is selecting high-information points to query and that it gives competitive performance compared to two other recent active learning methods: LAND and LEND. Specifically, MASC consistently outperforms these algorithms in terms of computation time and exhibits competitive accuracy on Indian Pines for a broad range of query budgets.

### v2

@@ 4 Background
where $\Phi_{n}$ is the kernel defined in (3.3). The localization property (3.5), used with $\lceil S/2\rceil$ in place of $S$, implies that
@@ 4.1 Measures
In the case of signal separation, we have seen that if the minimal amplitude for a certain point source is sufficiently small, we may not be able to detect that point source. Likewise, if the measure $\mu$ is too small on parts of $\mathbb{X}$}, we may not be able to detect those parts. For this reason, we make some assumptions on the measure $\mu$ as in [cloningercluster]. The first property, detectability, determines the rate of growth of the measure locally around each point in the support. The second property, fine-structure, relates the measure to the classification problem by equipping the support with some well-separated (except maybe for some subset of relatively small measure) partition which may correspond to some different class labels in the data.
@@
Supposing that $\mu=\sum_{k=1}^{K}a_{k}\delta_{\omega_{k}}$ as in the signal separation problem, then we see that $\mu$ is detectable with $\alpha=0$, $\kappa_{1}=\sum_{k}|a_{k}|$, $\kappa_{2}=\min_{k}|a_{k}|$. In this context, the value $\kappa_{2}$ is known as the minimum weight and plays an important role in signal recovery (recall the threshold $\min_{k}\left|a_{k}\right|/2$ from Example 3.1). The measure $\mu$ also has fine structure in the classical sense whenever $\eta<\min_{j\neq k}|\omega_{j}-\omega_{k}|$. In this sense, the theory presented in this paper is a generalization of results for signal separation in this regime. ∎
@@ 4.2 F-score
We will give results on the theoretical performance of our measure estimating procedure by giving an asymptotic result involving the so-called F-score. The F-score for binary classification (true/false) problems is a measure of classification accuracy taking the form of the harmonic mean between precision and recall. In a predictive model, precision is defined as the fraction of true positive outputs over all the positive outputs of the model. Recall is the fraction of true positive outputs over all the actual positives. In a multi-class problem, we extend this definition as follows (cf. [diagraphcluster_ohiostate_2011]). If $\{C_{1},\dots,C_{N}\}$ is a partition of $\{x_{j}\}_{j=1}^{M}$ indicating the predicted output labels of a model and $\{L_{1},\dots,L_{K}\}$ is the ground-truth partition of the data, then one can define the precision of $C_{j}$ against the true label $L_{k}$ by $|C_{j}\cap L_{k}|/|C_{j}|$ and the corresponding recall by $|C_{j}\cap L_{k}|/|L_{k}|$. Taking the maximum of the harmonic means of the precisions and recalls with respect to all the ground truth labels leads to
@@ 5 Main results
Intuitively, due to the localization (4.2) of $\Psi$, we expect $F_{n}$ to be large when nearby any of the data points $x_{j}$, and small otherwise. With this understanding, $\mathcal{G}_{n}$ then thresholds points in the metric space where $F_{n}$ is sufficiently large. This gives an estimation for the support of the data. The thresholding value $\Theta$ decides the cutoff based on the maximal value that $F_{n}$ attains over the data. If $\Theta$ is too large we would expect to get an underestimation of the support set $\mathbb{X}$, whereas if it is too small we will get an overestimation. We formalize this intuition in our first theorem.
Remark 5.1.
The function $F_{n}$ resembles a kernel density estimator, and it would be if the data measure $\mu$ was assumed to be absolutely continuous with respect to some base measure on the ambient metric space $\mathbb{M}$. However, we do not assume a base measure on the ambient metric space. Moreover, even if the metric space were the Euclidean space or a sphere as we have used in Section 7, we do not require $\mu$ to have a density with respect to the natural base measure on these spaces; indeed, we are interested in the cases when $\mu$ is a singular measure. Finally, it is not our goal to approximate the measure itself, but merely to approximate its support. ∎
@@
Let $\mu$ be detectable and suppose $M\gtrsim n^{\alpha}\log(n)$. Let $\{x_{1},x_{2},\dots,x_{M}\}$ be independent samples from $\mu$. There exists a constant $C>0$ such that if $\Theta<C<1$, then there exists $r(\Theta)\sim\Theta^{-1/(S-\alpha)}$ (recall $S$ is the localization parameter of the kernel, given in (3.5) and (4.2)) such that with probability at least $1-c_{1}/M^{c_{2}}$ we have
@@
If $\mathcal{C}=\{z_{1},\cdots,z_{M}\}$ is a random sample from $\mu$, $n\geq 1$ and $M\gtrsim n^{\alpha}\log n$, then it can be shown (cf. [mhaskar2020kernel, Lemma 7.1]) that for any point $x\in{\mathbb{X}}$, there exists some $z\in\mathcal{C}$ such that $\rho(x,z)\leq 1/n$. Hence, the Hausdorff distance between $\mathbb{X}$} and $\mathcal{C}$ is $\leq 1/n$. If $\mu$ has a fine structure in the classical sense, and $n\gtrsim\eta^{-1}$, then this implies that a correct clustering of $\mathcal{C}$ would give rise to a correct classification of every point in $\mathbb{X}$}. This justifies our decision to construct the algorithm in Section 6 to classify only the points in $\mathcal{C}$. On the other hand, the use of the localized kernel as in the theorems above guide us about the choice of the points at which to query the label. ∎
@@
Figure 4: Demonstration of the support estimation set $\mathcal{G}_{32}(0.15)$ (yellow) applied to a simple two-moons data set from [twomoons] (blue and red). By querying one point from each component of the support estimation set and extending the label to the other points in the same component, we can classify the entire data set with 100% accuracy.
@@ 6.1 Algorithm description
MASC will collect all points which do not yet have predicted labels (line 17), and apply the nearest-neighbors approach as described above to each of these points (lines 19 and 20). At this point, every element in the data set will have a predicted label, so the algorithm will return the list of labels (line 21).
@@ 6.2 Comparison with CAC and SCALe
In [cloningercluster], a similar theoretical approach to this paper except on the Euclidean space was developed and an algorithm we will call “Cautious Active Clustering” (CAC) was introduced. MASC and CAC are both multiscale algorithms using $\mathcal{G}_{n}(\Theta)$ to threshold the data set, then constructing graphs to query points and extend labels. The main difference between the algorithms is the following. In CAC, $\eta,\Theta$ are considered hyperparameters while $n$ is incremented, whereas in MASC, $n,\Theta$ are considered hyperparameters while $\eta$ is incremented. This adjustment serves three purposes:
@@
SCALe, as introduced in [mhaskar-odowd-tsoukanis] is an even more similar algorithm to MASC. The main difference between MASC and SCALe is the final step, where in the present method we use a nearest-neighbors approach to extend labels to uncertain points while in SCALe the choice was to use a function approximation technique developed in [mhaskarodowd]. Both methods have their pros and cons. Compared to SCALe, the nearest-neighbors approach of MASC:
@@
    works in arbitrary metric spaces, without requiring a summability kernel as in SCALe.
@@ 7 Numerical examples
In this section, we look at the performance of the MASC algorithm applied to 1) a synthetic data set with overlapping class supports (Section 7.1), 2) a document data set (Section 7.2), and 3) two different hyperspectral imaging data sets: Salinas (Section 7.3) and Indian Pines (Section 7.4). In each case, we project the (potentially pre-processed) data to the sphere and use $\rho=\arccos(\circ\cdot\circ)$ as the metric for graph construction. This guarantees that the metric space has diameter $\leq\pi$. On the Hyperspectral data sets, we compare our method with two other algorithms for active learning: LAND and LEND (Section 7.5).
For hyperparameter selection on our model as well as the comparisons, we have not done any validation but rather optimized the hyperparameters for each model on the data itself. So the results should be interpreted as being near-best-possible for the models applied to the data sets in question rather than a demonstration of generalization capabilities. While this approach is non-traditional for unsupervised/supervised learning, it has been done for other active learning research ([murphy-lend], for example) so we have elected to follow the same procedure in this paper. Further, an exhaustive grid search was not conducted but rather local minima among grid values were selected for each hyperparameter. For MASC, we looked at $n$ in powers of $2$ and $\overline{k}$ values in multiples of $5$. For LAND we looked at $K,t$ at increments of $10$, and with LEND we used the same parameters from LAND and looked at integer $J$ values and $\alpha$ values in increments of $0.1$. For $\Theta$, we tried values less rigorously, meaning that better $\Theta$ values may exist than the ones chosen. Due to the nature of the algorithm, increasing $\Theta$ will increase the number of samples that the nearest-neighbors approach has to estimate, while reducing the number of labeled neighbors it has to do so. However, increasing $\Theta$ can also reduce the number of queries used, sometimes without deterioration in accuracy. So there may be some tradeoff, but we generally see the best results when $\Theta$ is chosen to threshold a small portion of the initial data (outlier removal). In Table 2, we summarize the choice of parameters for each of the data sets in the subsequent sections.
@@ 7.2 Document data
This numerical example uses the document data set provided by Jensen Baxter through Kaggle [documentdata]. The data set contains 1000 documents total, 100 each belonging to a particular category from: business, entertainment, food, graphics, historical, medical, politics, space, sport, and technology. For prepossessing we run the data through the Python sklearn package’s TfidfVectorizer function to convert the documents into vectors of length 1684. Then we implement MASC.
@@
Figure 6: This figure illustrates the classification process undergone MASC on the document data set at two points. On the left, we see the classification of points before the $\overline{k}$-nearest neighbors extension. On the right, we see the result after $\overline{k}$-nearest neighbors extension. Figure dimensions are the directions associated with the largest 2 singular values of the data matrix.
@@ 7.3 Salinas hyperspectral data
This numerical example is done on a subset of the Salinas hyperspectral image data set from [hsidata]. The full data set is visualized in Figure 9. Our subset of the Salinas data set consists of 20034 data vectors of length 204 belonging to 10 classes of the 16 original classes. Specifically, we took half of the data points at random from each of the first 10 classes of the original data set. For preprocessing we ran PCA and kept the first 50 components. Then we implemented MASC.
Figure 9: Visualization of the full Salinas data set ground truth labels by geographic location. Image sourced from [hsidata].
@@
Figure 10: This figure illustrates the classification process undergone by MASC at two points on the Salinas hyperspectral data set. On the left, we see the classification of points before the $\overline{k}$-nearest neighbors extension. On the right, we see the result after $\overline{k}$-nearest neighbors extension. Dimensions correspond to the second and third directions of greatest variance according to the PCA decomposition.
@@ 7.4 Indian Pines hyperspectral data
This numerical example is done on a 5-class subset of the Indian Pines hyperspectral image data set from [hsidata]. The full data set is visualized in Figure 13. Our subset of the Indian Pines data set consists of 5971 data vectors of length 200 belonging to classes number 2,6,11,14,16 of the 16 original classes. For preprocessing we normalized each vector. Then we implement MASC.
Figure 13: Visualization of the full Indian Pines data set ground truth labels by geographic location. Image sourced from [hsidata].
@@
Lastly in Figure 17, we show how points are assigned estimated labels over several sequential iterations of the algorithm. This figure demonstrates the multiscale approach of the algorithm, wherein high density points tend to be assigned labels early on in the iterations and low density points tend to be assigned labels later. Black points in the figure correspond to those not yet given labels after the shown iterations. Some of the black points correspond to very low density points which may have been thresholded out at the beginning of the algorithm process, while others correspond to points of potential label conflict to be resolved after the iterations.
@@
Figure 14: This figure illustrates the classification process undergone by MASC at two points on the Salinas hyperspectral data set. On the left, we see the classification of points before the $\overline{k}$-nearest neighbors extension. On the right, we see the result after $\overline{k}$-nearest neighbors extension. Figure dimensions are the directions associated with the largest 2 singular values of the data matrix.
@@
Figure 17: Visualization of label assignment by iteration. Shown are the points given new estimated labels by the time MASC has reached the iteration with the shown $\eta$ value. Also depicted in the legend of the plot is the number of queries used by said iteration. The iterations are shown sequentially ($\eta$ increasing) but not consecutively (intermediate iterations not shown and new points from those iterations lumped into the shown iterations).
@@ 7.5 Comparison with LAND and LEND
We compare our method with the LAND [murphy-land] algorithm and its boosted variant, LEND [murphy-lend]. In Figure 18, we see the resulting accuracy that each algorithm achieves on both Salinas and Indian Pines for various query budgets. On the left, we observe that our method achieves a comparable accuracy to both LAND and LEND at around 50 queries, then gradually surpasses the accuracy of LAND as the number of queries surpasses around 200. On the right, our method achieves a lower accuracy for a small number of queries, but then outperforms both LAND and LEND after the budget exceeds about 60 queries.
@@ 8.1 Measure support estimation
In order to prove this lemma, we first recall a consequence of the Bernstein inequality for trigonometric polynomials ([paibk], Chapter III, Section 3, Theorems 1 and Lemma 5).
@@ 8.2 Discretization
Information and a derivation for the Bernstein concentration inequality are standard among many texts in probability; we list [concineq, Section 2.1, 2.7] as a reference. It is instinctive to use Proposition 8.1 with $X_{j}=\Psi_{n}(x,x_{j})$. This would yield the desired bound for any value of $x\in\mathbb{M}$. However, to get an estimate on the supremum norm of the difference $F_{n}-\sigma_{n}$, we need to find an appropriate net for $\mathbb{M}$ (and estimate its size) so that the point where this supremum is attained is within the right ball around one of the points of the net. Usually, this is done via a Bernstein inequality for the gradients of the objects involved. In the absence of any differentiability structure on $\mathbb{M}$, we need a more elaborate argument.
The following proposition is a consequence of [mhaskar2020kernel, Theorem 7.2], and asserts the existence of a partition of $\mathbb{X}$ satisfying properties which will be helpful to proving our main results.
@@
Remark 8.1.
We note that the proof of Lemma 8.4 involves only the upper bound of the detectability condition (Definition 4.1). The lower detectability bound instead plays a role in deducing the lower bound on $J_{n}$ (see (8.19)) which in turn aids in deducing lower bounds in Theorems 8.1 and 8.2.
@@ 9 Conclusions
We also give some empirical analysis for the performance of our new algorithm MASC, which was originally introduced in a varied form in [mhaskar-odowd-tsoukanis]. The key focus of the algorithm is on querying high-information points whose labels can be extended to others belonging to the same class with high probability. This is done in a multiscale manner, with the intention to be applied to data sets where the minimal separation between supports of the measures for different classes may be unknown or even zero. We applied MASC to a document data set as well as two hyperspectral data sets, namely subsets of the Indian Pines and Salinas hyperspectral imaging data sets. In the process of these experiments, we demonstrate empirically that MASC is selecting high-information points to query and that it gives competitive performance compared to two other recent active learning methods: LAND and LEND. Specifically, MASC consistently outperforms these algorithms in terms of computation time and exhibits competitive accuracy on Indian Pines for a broad range of query budgets.